In face centered cubic (fcc) crystal lattice, edge length is 400 pm. Find the diameter of greatest sphere which can be fit into the interstitial void without distortion of lattice.

Grade:upto college level

2 Answers

Deepak Patra
9 years ago
Hello Student,
Please find the answer to your question
For an octahedral void a = 2 (r + R) In fcc lattice the largest void present is octahedral void. If the radius of void sphere is R and of lattice sphere is r. Then,
r = √2 *400/5 = 141.12 pm (a = 400 pm)
applying condition for octahedral void, 2 (r + R) = a
∴ 2 R = a 2r = 400 – 2 * 141.12
∴ Diameter of greatest sphere = 117.16 pm

Thanks
Deepak patra
askIITians Faculty
Ayush Bhardwaj
6 years ago
In fcc lattice, thr largest void present is octahedral void and its radius(R) is given byR=(a-2r)/2Where,a= edge length,r=radius of the lattice atomBut r=?We know thatr=√2 × a/4r=1.414×400/4r=141.4pmNow diameter of the biggest sphere=2RFrom relation given above,2R=a - 2r 2R = 400 - 2(141.4) 2R = 400 - 282.6 2R = 117.4pmHence the diameter of the biggest sphere that can be inserted is 141.4pm